# Aristotle’s Second-Order Logic

### v8

@@ Conclusion
We now discuss in light of the present work the system of argumentation for which the Topics was written. In the Platonic Dialogues (particularly the early222222in a future paper we show the surprising second-order formalizability of the earlier Socratic dialogues as contrasted to the later ones. and middle ones) as well as in Aristotle there is the fundamental supposition that philosophical knowledge can be gained by finding the right definition of terms (under the proviso that the term in question is acknowledged to exist, that is, have non-empty extension.). This concern is not abandoned in the Analytics. The Socratic attitude is that the first step to wisdom consists in testing one’s beliefs which can lead to the valuable cathartic knowledge of the error of one’s beliefs. Thus a major part of philosophical activity involved the testing (i.e. attempt at refutation) of proposed definitions of terms - something which Aristotle acknowledges to be easier than finding correct definitions [^23] A massive amount of arguments were compiled during the eras of the late Academia and first Stoic schools, as testified by accounts of the lives of Chrysippus and Carneades.. In a modern framework how are we to understand (broadly speaking) the dialectical game of the Topics ? In ASOL in which we assume we have a finite collection of constants, functions and predicate symbols (first and second order) which represent the disambiguated vocabulary of a common language. The non-logical axioms we assume are $T$, the topics (or rather a minimal generating set for them ) and $K$ those representing common agreed upon knowledge. The game begins with Alice’s proposal of a definition of a certain term $hAD$. Then Bob basically has to try to derive a contradiction $T+K+hAD\vdash\bot$. The way the game was played apparently involved Alice confirming every move made by Bob, as attested in the Socratic dialogues (where it is rare for the ’opponent’ not to agree with Socrates’ moves). The ’moves’ can be $K$ axioms themselves or inferences, either purely logical or based on a topic. A nice example of this dialectical game in practice is given by Plato’s Lysis (see the ASOL formalization of a key section in appendix 2) although the types of definition considered do not fall into Aristotle’s scheme for monadic predicates terms outlined in book VI of the Topics. As discussed at the end of the previous section, it would be interesting to study the early and middle Platonic dialogues from a second-order perspective as in the present work so as to provide additional evidence for logical competency in dealing with multiple generality among the ancients.
@@
We can of course question if this dialectical game could lead to actual progress in philosophy or to certain or plausible knowledge about reality. In particular the exact scope of $K$ seems to be vague and uncertain; the assumption of the possession of a perfectly disambiguated language is itself implausible (although it should be said that the Topics is acutely aware of this problem). However we should not go to the other extreme of discarding this kind of philosophical practice. Rather what was required is a preliminary radical epistemic, semantic and logical analysis in order to set up a relevant and meaningful system - and such concerns are in no way foreign to the mature Aristotle: things better known to us are not necessarily better known in themselves242424as we saw in section 5, the extra-logic second-order relation of ’being better known’ plays an important role in the Topics: if $A$ is a property of $B$ then it has to be better known than $B$. In book I,1 Aristotle says that a syllogism is a ’proof’ when the premises are ’true and primitive’ or epistemically depending on ’true and primitive’ premises. The adjective ’alethê’ is also translated as ’evident’. Aristotle then clarifies that evident and primitive is what is known through itself and not through another’. All ’scientific principles’ should be self-evident. Other propositions are ’endoksa’, translated sometimes as ’probable’..
@@
9. $\sim\forall xSx(\iota yR^{\ast}xy)\rightarrow\sim\mathfrak{g}RS$ (124b)252525if $R$ is one-to-one.
@@
Appendix 4: the rules of ASOL
Prawitz[12] gives two versions of second-order logic (SOL) and ASOL can be built upon either of these versions. Here we only consider version 2 and adapt the notation to the use of the present paper. We define formula, individual term and $n$-ary predicate term as usual in second-order logic by simultaneous recursion as in [12]. Individual variables (which are individual terms) are denoted by $x,y,z,..$ and $1$-ary (monadic) predicate term variables (which are $1$-ary predicate terms) are denoted by $A,B,C,...S,T,...$. We denote a $n$-ary predicate term variable by $P^{n},X^{n},...$. Given a formula $\phi$ we can also form the $n$-ary predicate term $\lambda x_{1}....x_{n}\phi$. In the rules ahead $t$ denotes a variable term and $\mathfrak{T}^{n},\mathfrak{P}^{n}$ a $n$-ary predicate term. The language of ASOL is obtained by adding to SOL second-order predicate symbols $\alpha,\beta,...$ and second-order functions $f,g,...$. Each symbol is equipped with a signature which specifies its number of arguments (at least one) and for each argument whether it is an individual argument or a predicate argument, in which case the arity is also specified. For example $\gamma XY$ of section 4 is specified by having two predicate arguments. Note that if the arguments are all individual then the second-order predicate can be identified with a predicate term constant.
@@
where in $\forall^{1}I$ $a$ must not occur free in any assumption on which $\phi$ depends and in $\exists^{1}I$ $a$ must not occur free in $\psi$ or in $\exists x\phi$ or on any hypothesis on which $\psi$ depends except for $\phi^{x}_{a}$. Here $\phi^{t}_{s}$ denotes the result of substituting $t$ by $s$ in $\phi$ [^26] Here we need not follow Prawitz’s terminological distinction between variables and parameters which correspond to bound and free variables..
@@
Finally there are introduction and elimination rule for $\lambda$-predicate terms concerning which the reader may consult [12].

### v11

@@ Conclusion
We now discuss in light of the present work the system of argumentation for which the Topics was written. In the Platonic Dialogues (particularly the early262626in a future paper we show the surprising second-order formalizability of the earlier Socratic dialogues as contrasted to the later ones. and middle ones) as well as in Aristotle there is the fundamental supposition that philosophical knowledge can be gained by finding the right definition of terms (under the proviso that the term in question is acknowledged to exist, that is, have non-empty extension.). This concern is not abandoned in the Analytics. The Socratic attitude is that the first step to wisdom consists in testing one’s beliefs which can lead to the valuable cathartic knowledge of the error of one’s beliefs. Thus a major part of philosophical activity involved the testing (i.e. attempt at refutation) of proposed definitions of terms - something which Aristotle acknowledges to be easier than finding correct definitions [^27] A massive amount of arguments were compiled during the eras of the late Academia and first Stoic schools, as testified by accounts of the lives of Chrysippus and Carneades.. In a modern framework how are we to understand (broadly speaking) the dialectical game of the Topics ? In ASOL in which we assume we have a finite collection of constants, functions and predicate symbols (first and second order) which represent the disambiguated vocabulary of a common language. The non-logical axioms we assume are $T$, the topics (or rather a minimal generating set for them ) and $K$ those representing common agreed upon knowledge. The game begins with Alice’s proposal of a definition of a certain term $hAD$. Then Bob basically has to try to derive a contradiction $T+K+hAD\vdash\bot$. The way the game was played apparently involved Alice confirming every move made by Bob, as attested in the Socratic dialogues (where it is rare for the ’opponent’ not to agree with Socrates’ moves). The ’moves’ can be $K$ axioms themselves or inferences, either purely logical or based on a topic. A nice example of this dialectical game in practice is given by Plato’s Lysis (see the ASOL formalization of a key section in appendix 2) although the types of definition considered do not fall into Aristotle’s scheme for monadic predicates terms outlined in book VI of the Topics. As discussed at the end of the previous section, it would be interesting to study the early and middle Platonic dialogues from a second-order perspective as in the present work so as to provide additional evidence for logical competency in dealing with multiple generality among the ancients.
@@
We can of course question if this dialectical game could lead to actual progress in philosophy or to certain or plausible knowledge about reality. In particular the exact scope of $K$ seems to be vague and uncertain; the assumption of the possession of a perfectly disambiguated language is itself implausible (although it should be said that the Topics is acutely aware of this problem). However we should not go to the other extreme of discarding this kind of philosophical practice. Rather what was required is a preliminary radical epistemic, semantic and logical analysis in order to set up a relevant and meaningful system - and such concerns are in no way foreign to the mature Aristotle: things better known to us are not necessarily better known in themselves282828as we saw in section 5, the extra-logic second-order relation of ’being better known’ plays an important role in the Topics: if $A$ is a property of $B$ then it has to be better known than $B$. In book I,1 Aristotle says that a syllogism is a ’proof’ when the premises are ’true and primitive’ or epistemically depending on ’true and primitive’ premises. The adjective ’alethê’ is also translated as ’evident’. Aristotle then clarifies that evident and primitive is what is known through itself and not through another’. All ’scientific principles’ should be self-evident. Other propositions are ’endoksa’, translated sometimes as ’probable’..
@@
9. $\sim\forall xSx(\iota yR^{\ast}xy)\rightarrow\sim\mathfrak{g}RS$ (124b)292929if $R$ is one-to-one.
@@
Appendix 4: Presentation of ASOL and ESOL
ASOL is a special case of extended second-order logic (ESOL) defined as follows. We are given a collection of individual variables $x,y,z,...$ and individual constants $a,b,c,...$ together with a collection of $n$-ary predicate term variables $X^{(n)},Y^{(n)},Z^{(n)},...$ (we omit the superscript if the context is clear) for $n\geq 0$. Furthermore we are given a collection of predicate term symbols $\alpha^{(\tau)},\beta^{(\tau)},\gamma^{(\tau)},...$ where $\tau$ consists of a (possibly empty) list $(n_{1},...,n_{m})$ of integers $n_{i}\geq-1$ called a signature, where we likewise omit the superscript if the context is clear. It is also convenient to consider that we have a collection of function symbols $f^{(\tau,n)},g^{(\tau,n)},...$ where $\tau$ is a signature and $n\geq-1$. For $n=-1$ these are called first-order functional symbols. Otherwise they are called second-order function symbols.
We define formula and $n$-predicate term and individual term by mutual recursion.
For convenience we use small-case Roman letters to represent individual terms, variables and constants and upper-case Roman letters to represent $n$-predicate terms and variables. We use gothic letters to represent terms in general. We use upper-case Greek letters to represent formulas.
If $\Phi_{1},\Phi_{2}$ are formulas then so are $(\Phi_{1}\vee\Phi_{2}),(\Phi_{1}\wedge\Phi_{2}),\neg\Phi_{1},\forall x\Phi_{1},\exists x\Phi_{1},\forall X\Phi_{1},\exists X\Phi_{1}$ where $x$ is an individual variable and $X$ is a predicate term variable.
If $\alpha^{(\tau)}$ is a predicate term symbol with $\tau=(n_{1},...,n_{m})$ non-empty then $\alpha\mathfrak{t}_{1}....\mathfrak{t}_{m}$ is a formula where if $n_{i}\geq 0$ then $\mathfrak{t}_{i}$ is a $n_{i}$-predicate term and if $n_{i}=-1$ then $\mathfrak{t}_{i}$ is an individual term. If $\tau$ is empty then $\alpha$ is considered a 00-predicate term.
If $\Phi$ is a formula and $x_{1},...,x_{n}$ is a list of individual variables (possible empty) then $\lambda x_{1}...x_{n}\Phi$ is a $n$-predicate term (for the empty list we put $n=0$ and write $\lambda\Phi$.
If $X^{(n)}$ is a predicate term variable then it is a $n$-predicate term. If $\alpha^{(\tau)}$ is a predicate term symbol such that $\tau=(k_{1},...,k_{n})$ with $k_{i}=-1$ then $\alpha^{(\tau)}$ is a $n$-predicate term303030we can also call $\alpha$ in this case a $n$-predicate constant.
If $T$ is a $n$-predicate term and $t_{1},...t_{n}$ are individual terms then $Tt_{1}...t_{n}$ is a formula. For the case $n=0$ we write the formula $T()$.
An individual constant or variable is an individual term.
If $f^{(\tau,n)}$ is a function symbol then $f\mathfrak{t}_{1}...\mathfrak{t}_{m}$ is an individual term (for $n=-1$) or $n$-predicate term (for $n\geq 0$) where $\mathfrak{t}_{1}...\mathfrak{t}_{m}$ satisfy the condition relative to $\tau$ as above.
The natural deduction rules consist of the standard rules for the first-order classical predicate calculus:
@@
where in $\forall^{1}I$ $y$ must not occur free in any assumption on which $\Phi$ depends and in $\exists^{1}I$ $y$ must not occur free in $\Psi$ or in $\exists x\Phi$ or on any hypothesis on which $\Psi$ depends except for $\Phi^{x}_{y}$. Here $\Phi^{t}_{s}$ denotes the result of substituting $t$ by $s$ in $\Phi$ [^31] Here we need not follow Prawitz’s terminological distinction between variables and parameters which correspond to bound and free variables..
@@
Finally there are introduction and elimination rule for $\lambda$-predicate terms for which the reader may consult [15].
In Prawitz[15] versions 1 and 2 are proof-theoretically equivalent (for formulas without $\lambda$), the difference being that in version 1 the rules $\lambda I$ and $\lambda E$ are incorporated into second-order quantifier rules.
Why do we consider ESOL as being an extension of second-order logic rather than a fragment of third-order logic ? We are of the view that what defines the ’order’ of a logic is what is quantified over and what kind of corresponding variable there is. First-order logic quantifies over individuals and has individual variables. Second-order logic quantifiers over entities corresponding to $n$-ary predicates with individual arguments and has the corresponding $n$-ary predicate variables. Third-order logic thus would have to quantify over entities corresponding to predicates which can involve both individual arguments and $n$-ary predicates of individuals arguments and furthermore it would require variables to represent these last. Such a quantification and such variables are lacking in ESOL although the predicate term symbols $\alpha^{(\tau)}$ can be considered in some sense as being ’third-order’ when involving $n$-predicate term arguments.
There are some extensions and modifications of ESOL which can make it even more adequate to formalize Aristotle’s reasoning. These also are interesting in their own right. Thus our project will involve the following:
  1. 1.
[…]
    The reason for this investigation is to develop my alternative proposal for an interpretation of Aristotle’s demonstrative technique of ekthesis which calls for a generalized indefinite version of the Peano iota operator (while the Peano operator corresponds to a definite article this construction corresponds to the indefinite article). If we have that there exists an A, then we can form the term ’an A’ so as to be able to proceed thus: let a be an A. An example of a natural language expression of the rule above: every man has a father, hence, every man’s father is the father of that man (or: if every man has a father then a father of a man is that man’s father).
  2. 2.
    In natural language discourse most (if not all) quantifiers appear to be bounded. This is true of Aristotle’s reasoning and this approach has proven itself useful in contemporary applications of dependent type theory. Thus we will investigate alternative forms of the natural deduction rules for ESOL in which all quantifiers are bounded $\forall x(\Phi(x)\rightarrow\Psi(x))$ (we write this $\forall_{x\Phi}\Psi$). In this alternative natural deduction system Modus Ponens is incorporated into the quantifier elimination rules, much as the $\lambda$ rules are incorporated into the quantifier rules in Prawitz’s first version of second-order logic. Some of the rules might look like:
[…]
    Rule $\forall^{1}E$ is of course the famous syllogism involving singular propositions (not in fact found in Aristotle’s Organon): all men are mortal, Socrates is mortal, therefore Socrates is mortal. To bring the logic closer to natural language we could also construct an alternative system in which all formulas (including the formulas appearing in the rules) are closed formulas. This would require for instance Modus Ponens being incorporated into the new quantifier laws or vice-versa. For instance we would have to replace MP we rules like:
[…]
    If we combine this with bounded quantifiers then it seems we would require instead of MP extensions like:
[…]
    which is just the old Barbara syllogism.
  3. 3.
    Recall that starting from $\forall^{2},\forall^{1},\rightarrow$ we can define all the other connectives. We can define ESOLat in analogy to atomic polymorphism for system $F$ by restricting $\forall E$ to predicate terms without $\lambda$. We can embed system $F$ ($\lambda 2$) into ESOL. Furthermore this embedding can be used to show that ESOL is undecidable due to $F$ being undecidable. Does the undecidability of $F_{at}$[1] carry over to ESOLat ?
  4. 4.
    There is a certain redundancy in our definition of formulas and terms in ESOL. In the definition of formula in ESOL the construction of a formula via the application of an $n$-predicate term to a list of individual variables $Xt_{1}...t_{n}$ can be considered as a special case of the application of predicate symbols (see the previous definition) via a special predicate term symbol $\Delta^{n}$ with signature $(n,-1,...,-1)$ satisfying axiom:
    $Xt_{1}....t_{n}\leftrightarrow\Delta^{n}Xt_{1}...t_{n}$
    We propose to explore this alternative presentation of ESOL. Another topic is to clarify the relative strength and relationship between ESOL and multi-sorted first-order logic as well as Bealer’s intensional logic (such as presented, for instance, in Quality and Concept[2]).
On a philosophical side we plan to develop the following arguments for the philosophical interest as well as applicability of ESOL.
  1. 1.
    One reason to consider ESOL as a philosophical logic arises when we are confronted with situations in which we wish to distinguish intention from extension, the meaning of a concept from its extension. For example for Aristotle the concept ’able to laugh’ and ’rational animal’ both have humans as their extension but their meanings are clearly different. It is necessary to have an equality predicate defined not only for individuals but for $n$-ary predicate terms, in particular monadic predicate term constants. This equality $C^{1}=D^{1}$ need not be equivalent to $\forall x(Cx\leftrightarrow Dx)$.
  2. 2.
    Zalta’s Object Logic [25], which in one presentation is a second-order logic with fusions of concepts into abstract objects, provides an alternative interpretation of Fregean Arithmetic. It is based on a primitive encoding relation $enc\,tT^{1}$ which involves an individual term $t$ and 1-ary predicate term $T^{1}$. Object Logic is thus a particular case of ESOL.
  3. 3.
    In order to formalize propositional attitudes it is convenient to have predicates which can take 00-ary predicate terms as arguments. ESOL meets this need. Note that in ESOL a 00-ary predicate term $T^{0}$, for instance $\lambda\Phi$ for some sentence $\Phi$ can be construed as expressing the meaning of $\Phi$ while $T^{0}()$ expresses the judgment involving a truth-value claim. Thus the relation $K$ of agent $a$ knowing sentence $\Phi$ would be expressed as $Kx\lambda(\lambda\Phi)())$.
  4. 4.
    In certain situations (as in ’Zig-Zag and Fregean Arithmetic’[8]) we want to distinguish between predicative and non-predicative predicates, concepts which are ’too complicated’ to have extensions (analogously to how some extensions are ’too complicated’ to be the extension of a single concept). This again calls for a monadic predicate having predicate term arguments, for instance $\pi X^{(1)}$ expressing that $X^{(1)}$ is predicative.
  5. 5.
    We can also consider functions taking predicate terms as arguments and forming either individual terms or predicate terms. The later case could be used to interpret first and second-order multi-modal logic as already suggested in this paper.